OpenStudy (anonymous):

CAN ANY ONE TELL ME THE FORMULA TO DETERMINE THE STANDARD VALUE OF AN ITEM , WHEN DISCOUNT VALUE AND PRICE OF ITEM AFTER DISCOUNT IS GIVEN ?

OpenStudy (anonymous):

STANDARD VALUE=x DISCOUNT VALUE=y PRICE OF ITEM AFTER DISCOUNT=z z=x-y x=z+y

OpenStudy (anonymous):

DISCOUNT VALUE IS GIVEN IN PERCENTAGE FORM AND THE STANDARD PRICE AND THE PRICE AFTER DISCOUNT ARE IN CURRENCY FORM